Consider a sample with a mean of 40 and a standard deviation of 5. Use Chebyshev's theorem to determine the percentage of the data within each of the following ranges (to the nearest whole number).

  1. 20 to 60, at least ? %
  2. 15 to 65, at least ? %
  3. 31 to 49, at least ? %
  4. 27 to 53, at least ? %
  5. 24 to 56, at least ? %
Answer #1


Formulas: 1) If ơ is standard deviation of the random variable X then the variance of random variable X is denoted by var(X) Now we have to find the percentage of the given data within each of the following ranges using chebychevs theorem. a) P(20 Xb) P(15 X 65) = P(15-40 X-E(X) 65-40) = P(-25 X-E(X) 25) -PCIX-E(X)| 25) var(X) 21- _25 (25)2 24 2 25 2 0.96 P(15 X 65) = atlc) P(31 X 49) = P(31-40 X-E(X) 49-40) var(X) 21 25 (9)2 56 81 0.6914 15 X 65) = atleast 69.14% = 69%d) P(27 X 53) = P(27-40 X-E(X) 53-40) - P(-13 sX - E(X) s 13) -PCIX-E(X)| 13) var(x) 21- _25 144 169 2 0.8521 (13)2 P(15 X 65